# Delin Qu∗1, Haoming Song∗1, Qizhi Chen∗1, Yuanqi Yao1, Xinyi Ye1, Yan Ding1, Zhigang Wang1 Jiayuan Gu3, Bin Zhao1, Dong Wang1, Xuelong Li12 1Shanghai AI Laboratory, 2TeleAI, 3ShanghaiTech

### v4

@@ Abstract
In this paper, we claim that spatial understanding is the keypoint in robot manipulation, and propose SpatialVLA to explore effective spatial representations for the robot foundation model. Specifically, we introduce Ego3D Position Encoding to inject 3D information into the input observations of the visual-language-action model, and propose Adaptive Action Grids to represent spatial robot movement actions with adaptive discretized action grids, facilitating learning generalizable and transferrable spatial action knowledge for cross-robot control. SpatialVLA is first pre-trained on top of a vision-language model with 1.1 Million real-world robot episodes, to learn a generalist manipulation policy across multiple robot environments and tasks. After pre-training, SpatialVLA is directly applied to perform numerous tasks in a zero-shot manner. The superior results in both simulation and real-world robots demonstrate its advantage of inferring complex robot motion trajectories and its strong in-domain multi-task generalization ability. We further show the proposed Adaptive Action Grids offer a new and effective way to fine-tune the pre-trained SpatialVLA model for new simulation and real-world setups, where the pre-learned action grids are re-discretized to capture robot-specific spatial action movements of new setups. The superior results from extensive evaluations demonstrate the exceptional in-distribution generalization and out-of-distribution adaptation capability, highlighting the crucial benefit of the proposed spatial-aware representations for generalist robot policy learning. All the details and codes will be open-sourced.
@@ I Introduction
Generalist robot policies that are capable of interacting with the physical environment, adapting to various embodiments, and performing complex tasks have been a long-standing pursuit in robotics [6, 3, 16, 8, 60]. Recent advances in Vision-Language-Action (VLA) models [7, 29, 5, 32] show a promising paradigm in building such generalist policy by fine-tuning the pre-trained Vision-Language Models (VLMs) [1, 51, 48, 36] on diverse robot data [13, 28, 18]. The key to the success of this paradigm lies in adapting the generalization power of VLMs to numerous robot manipulation tasks, as well as specific architectural designs that synergize the VLM backbone and robot action output head. Nonetheless, existing VLA models are primarily confined to 2D observation inputs and lack precise perception and comprehension of the 3D physical world — where humans instinctively construct rich, structured mental representations of space, effortlessly aligning objects within a canonical, intuitive, and even personally tailored workspace for manipulation. Therefore, an essential question for the field now is how to effectively equip the VLA models with a profound spatial understanding of the 3D physical world?
@@ II Related Work
3D Foundation Models for Robotics. Some researches [67, 10, 19, 23, 24, 63] have focused on extending the generalist ability of LLMs and VLMs from language-vision towards the 3D world. 3D-LLM [23] integrates a 3D feature extractor with 2D VLMs backbone and train 3D-LLMs on a wide variety of tasks, including dense captioning, 3D question answering, task decomposition, 3D grounding, 3D-assisted dialog, navigation, and so on. LLaVA-3D [10] extends the 2D LLaVA’s capabilities with the proposed 3D patches to bridge 2D features within a 3D space for 3D spatial understanding. Similarly, LEO [24] trains an embodied multi-modal generalist agent that can take egocentric 2D images, 3D point clouds, and texts as task input and handle comprehensive tasks within the 3D environment. Moreover, 3D-VLA [63] builds a generative world model on top of 3D-based LLM to perform 3D reasoning and localization, multimodal goal generation, and embodied action planning. LEO and 3D-VLA are closely related to our work, but their attention is on 3D world understanding and prediction, ignoring the 3D spatial characteristic in the robot action space.
@@ III Methodology
In this section, we describe SpatialVLA model and its training framework in detail. Our SpatialVLA model with the proposed Ego3D position encoding and adaptive action grids to capture and learn 3D spatial knowledge for generalizable robot control, which we describe in Section. III-A. Next, we detail the training procedure of SpatialVLA that consists of a pre-training stage and a post-training stage in Section. III-B. The pre-training aims to learn generalizable knowledge with large-scale cross-robot data and the goal of post-training is to adapt pre-trained model to specific downstream robot embodiments and tasks.
@@ III-A The SpatialVLA Model Architecture
As illustrated in Figure. 1, SpatialVLA is developed based on a vision-language model to inherit the general world knowledge. Formally, SpatialVLA takes image observations $\mathbf{o}_{t}=\{\mathbf{I}_{t}^{1},..,\mathbf{I}_{t}^{n}\}$ and a natural language task instruction $\mathbf{L}$ as inputs, and then learns a mapping function $\tau(\cdot)$ to generate a sequence of robot actions $\mathbf{A}_{t}=[\mathbf{a}_{t},\mathbf{a}_{t+1},...,\mathbf{a}_{t+H-1}]$, i.e., $\mathbf{A}_{t}=\mathcal{F}(\mathbf{o}_{t},L)$. To empower SpatialVLA with 3D spatial intelligence, we augment the VLM backbone with robotics-specific 3D-aware inputs and outputs, namely, Ego3D Position Encoding and Adaptive Action Grids. The ego3D position encoding representation $\mathbf{O_{3d}}$ aims to capture 3D scene structure via integrating 3D spatial information with 2D semantic features. The adaptive action grids are designed to represent the continuous distribution of robot actions $\mathbf{a}$ with a set of discrete spatial action tokens $\mathfrak{a}=\{\mathfrak{a}^{1},...,\mathfrak{a}^{V}\}$. During training, SpatialVLA model is trained to take the ego3D position encoding representation $\mathbf{O_{3d}}$ and natural language task instruction $\mathbf{L}$ as inputs, and autoregressively generate spatial action tokens $\tilde{\mathfrak{a}}_{t}$ using the cross-entropy objective $\mathcal{L}$,
@@
where the predicted action tokens $\tilde{\mathfrak{a}}_{t}=\tau(\mathbf{O_{3d}},\mathbf{L})$ is the de-tokenized into continuous action signals $\mathbf{a}_{t}$ for robot control. More details of the model architecture and action encoding can be found in Section. -B.
Ego3D Position Encoding. The proposed Ego3D position encoding integrates depth information from the camera frame and image pixels to construct an egocentric 3D coordinate system, which eliminates the need for robot-camera extrinsic calibration and is agnostic to specific robot setups. Specifically, we use ZoeDepth [4] to estimate depth map $D$ and obtain pixel’s 3D position $\mathbf{p}=\{x,y,z\}$ in the egocentric 3D coordinate system via back-projection $\pi^{-1}$ with camera intrinsic parameters. Then, as illustrated in Figure. 1, we first employ SigLIP [62] visual encoder to extract 2D semantic visual features $\mathbf{X}\in\mathbb{R}^{d\times h\times w}$ to inherit the alignment between vision and language, and calculate their corresponding 3D positions $\mathbf{P}\in\mathbb{R}^{3\times h\times w}$ in the egocentric 3D coordinate system. The egocentric 3D positions $\mathbf{P}$ are then encoded into 3D position embeddings $\mathbf{P}^{{}^{\prime}}\in\mathbb{R}^{d\times h\times w}$ through a sinusoidal function $\boldsymbol{\gamma}(\cdot)$ following by a learnable MLP. The egocentric 3D spatial representations $\mathbf{O_{3d}}\in\mathbb{R}^{d\times h\times w}$ are obtained by adding 3D position embedding $\mathbf{P}^{{}^{\prime}}$ and 2D path visual tokens $\mathbf{X}$, depicted as follows,
@@
As illustrated in Figure. 2, for tokenizing continuous translation and rotation movements, we first normalize each action variable into $[-1,1]$ for each robot environment and statistic the translation and rotation movements $\Delta\mathbf{R}=\{\text{roll, pitch, yaw}\},\;\Delta\mathbf{T}=\{\text{$\phi,\theta,r$}\}$ on the whole dataset mixture (see Section. -F), following with a parameterized Gaussian distribution fitting $\mathcal{N}(\mu^{a},\Sigma^{a})$. Then, the continuous actions are split into M intervals $\mathbf{G}_{i=1,..,\text{M}}=\{[a_{1}=-1,a_{2}),...,[a_{\text{M-1}},a_{\text{M}}=1]\}$ with equal probability $1/\text{M}$ on each normalized action variable, i.e.,
@@
where $\mathbf{E}_{\text{trans}}\in\mathbb{R}^{d\times\text{M}_{\text{trans}}}$,   $\mathbf{E}_{\text{rot}}\in\mathbb{R}^{d\times\text{M}_{\text{rot}}}$,   $\mathbf{E}_{\text{grip}}\in\mathbb{R}^{d\times 2}$ denote the translation, rotation, and gripper actions, and the total number of action tokens is $V=\text{M}_{\text{trans}}+\text{M}_{\text{rot}}+2$. After training, these learned spatial action tokens capture general robot action knowledge and show a surprising ability in new robot embodiment adaption, as discussed in Section. III-B. Moreover, it is worth noting that the model only needs to generate 3 tokens for one-step robot actions rather than 7 tokens as in RT-1 [6], RT-2 [7] and OpenVLA [29], achieving in fast model inference speed.
@@ III-B The Pre-training and Post-training Scheme
Pre-training Procedure. We train SpatialVLA from Paligemma2 backbone [58] on a cross-robot dataset mixture with 1.1 Million real robot demonstrations $\{\zeta_{1},...,\zeta_{n}\}$, covering a diverse range of robot embodiments, scenes, and tasks. This pre-training dataset mixture consists of a subset of OXE [13] and the RH20T dataset [18] and we modify the mixture weights from OpenVLA [29] according to the real-word testing performance of individual dataset, which are exhibited in Section. -A. At the beginning of pre-training, the embeddings $\mathbf{E}_{\mathfrak{a}}$ of spatial action tokens and parameters of MLP in egocentric 3D spatial representation are randomly initialized, and then they are optimized during training, as well as the parameters of vision encoder and LLM backbone. At each training step, a batch of data pairs is extracted at random timesteps $t_{1},...,t_{B}$ from shuffled demonstrations $\{\zeta_{i},...,\zeta_{j}\}$, i.e., a batch of tuple $[(\mathbf{o}_{t_{1}},\mathbf{A}_{t_{1}},\mathbf{L}_{i}),...,(\mathbf{o}_{t_{B}},\mathbf{A}_{t_{B}},\mathbf{L}_{j})]$, and then SpatialVLA is trained with a standard auto-regressive next-token prediction objective in eq. 1. Importantly, the embeddings of text tokens $\mathbf{E}_{\text{text}}$ are frozen to maintain the general world knowledge in pre-trained VLM, and the experimental results show this frozen operation is beneficial for the instruction following ability. Moreover, as discussed in OpenVLA [29], DROID dataset [28] are removed from the data mixture for the final third of pre-training to improve the quality of the pre-trained SpatialVLA model.
@@ IV Experiment
To answer these questions, as shown in Figure. 3, we evaluate SpatialVLA’s capabilities across a representative spectrum of 7 different robot learning scenarios with 24 real-robot tasks and 3 simulation environments. Firstly, we evaluate SpatialVLA in both SimplerEnv [34] simulation and the real-world WidowX robot platform (BridgeV2 [59] [59] setups), testing its out-of-the-box control capabilities on different robots with setups matching the pre-training dataset. Second, we assess the fine-tuning efficacy of our method in both simulation and real-world settings, including LIBERO [35] and new Franka robot setups, to adapt to new robot environments and tasks. Then, we design 4 special tasks that require precise spatial understanding in 2 different real-world robot environments to test the effectiveness of spatial representations of SpatialVLA. Finally, we conduct comprehensive ablation studies on a mixture of Fractal [6] and BridgeV2 [59] datasets to verify our design decisions in SpatialVLA. For more details on evaluation setups, see Section. -D.
Implementation Details. The SpatialVLA model is pre-trained with 1.1 Million real-robot demonstrations from the OXE  [13] and RH20T dataset [18] on a cluster of 64 A100 GPUs for 10 days, using a batch size of 2048. For input robot observation, the SpatialVLA policy is only conditioned on one third-person camera and takes one image for constructing egocentric 3D spatial representations. For output robot actions, the SpatialVLA policy predicts a chunk of $T=4$ future actions (12 spatial action tokens from total $V=8194$ tokens) and executes the ensemble actions before predicting the next chunk. During inference, SpatialVLA requires 8.5GB of GPU memory and runs at approximately 20Hz on one NVIDIA RTX 4090 GPU to run evaluations in both simulation and real-world. For more details about model training and deployment, please refer to the Section. -C.
@@ IV-A Performing Zero-shot Robot Control
For a more comprehensive evaluation, we conduct experiments on a real-world WidowX robot platform from the BridgeData V2 evaluation [59]. As shown in Figure. 4, we design seven task suites for the WidowX robot, encompassing language grounding, semantic understanding (unseen background and poses), and motion distractors (manually move the object). All generalist manipulation policies, including Octo, RT-1-X, OpenVLA, and RoboVLM, are evaluated across 7 task suites with 11 trials each, resulting in a total of 77 rollouts. A more detailed breakdown of all tasks and policy settings can be found in the Section. -D.
SimplerEnv Evaluation of Google Robot and WidowX. Table. I summarizes the zero-shot and fine-tuning results across different manipulation policies on the Google Robot setup. On average, SpatialVLA achieves the highest overall visual matching and variant aggregation performance with a significant margin. Our SpatialVLA model yields 71.9% and 75.1% Visual Matching scores in zero-shot and fine-tuning settings, surpassing the second-best policy RoboVLM by +15.6% and +11.7% margins. Notably, our model trained from scratch on OXE mixture with RH20T surpasses the state-of-the-art closed-source model RT-2-X [13], achieving superior performance in Visual Matching (71.9% vs 60.7%) and Variant Aggregation (68.8% vs 64.3%), while using significantly fewer model parameters (3.5B vs 55B). Qualitatively, we find that SpatialVLA exhibits greater generalizability and robustness across diverse robotic manipulation tasks and environmental conditions, characterized by varying visual appearances, which is further supported by its superior performance in variant aggregation. Table. II summarizes the results across different manipulation policies on the WidowX setup. Our model surpasses the state-of-the-art RoboVLM policy, achieving overall success rates of 34.4% and 42.7%. Fine-tuning on the BridgeV2 Dataset yields a remarkable 100% success rate in the ”Put Eggplant in Yellow Basket” task, demonstrating the model’s exceptional zero-shot manipulation capability.
@@
Real-world WidowX Evaluation. Figure. 4 presents the results of the real-world ”out-of-the-box” evaluation in WidowX robot platform. We observe that, in simple single-task scenarios (#1 close microwave), all the policies exhibit some generalizability, successfully completing tasks in unseen environments. However, in moderately complex tasks (#3-7), most policies, such as RT-1-X, Octo, and RoboVLM struggle with manipulation, frequently encountering issues like object misidentification and grasp failures. Compared to OpenVLA, our method demonstrates superior robustness in handling motion disturbances (human-induced dynamic object movement in tasks #3 and #4), successfully tracking and grasping carrot and eggplant. Furthermore, in the instruction-following tasks (#5-7), our method demonstrates strong instruction-following ability, accurately executing tasks like picking up a green cup and placing it on a white plate, not a pink one, based on color descriptions in the prompts, outperforming OpenVLA and other generalist policies. Overall, SpatialVLA achieves a higher average success rate, showcasing robust and generalizable operation capabilities in unseen scenarios, objects, language grounding, and dynamic motions.

### v5

# Zhigang Wang1 Jiayuan Gu5, Bin Zhao†1,6, Xuelong Li1,6

@@ Abstract
In this paper, we claim that spatial understanding is the keypoint in robot manipulation, and propose SpatialVLA to explore effective spatial representations for the robot foundation model. Specifically, we introduce Ego3D Position Encoding to inject 3D information into the input observations of the visual-language-action model, and propose Adaptive Action Grids to represent spatial robot movement actions with adaptive discretized action grids, facilitating learning generalizable and transferrable spatial action knowledge for cross-robot control. SpatialVLA is first pre-trained on top of a vision-language model with 1.1 Million real-world robot episodes, to learn a generalist manipulation policy across multiple robot environments and tasks. After pre-training, SpatialVLA is directly applied to perform numerous tasks in a zero-shot manner. The superior results in both simulation and real-world robots demonstrate its advantage of inferring complex robot motion trajectories and its strong in-domain multi-task generalization ability. We further show the proposed Adaptive Action Grids offer a new and effective way to fine-tune the pre-trained SpatialVLA model for new simulation and real-world setups, where the pre-learned action grids are re-discretized to capture robot-specific spatial action movements of new setups. The superior results from extensive evaluations demonstrate the exceptional in-distribution generalization and out-of-distribution adaptation capability, highlighting the crucial benefit of the proposed spatial-aware representations for generalist robot policy learning. All the details and codes are open-sourced.
@@ I Introduction
Generalist robot policies that are capable of interacting with the physical environment, adapting to various embodiments, and performing complex tasks have been a long-standing pursuit in robotics [6, 3, 16, 8, 65]. Recent advances in Vision-Language-Action (VLA) models [7, 30, 5, 33] show a promising paradigm in building such generalist policy by fine-tuning the pre-trained Vision-Language Models (VLMs) [1, 55, 50, 37] on diverse robot data [13, 29, 18]. The key to the success of this paradigm lies in adapting the generalization power of VLMs to numerous robot manipulation tasks, as well as specific architectural designs that synergize the VLM backbone and robot action output head. Nonetheless, existing VLA models are primarily confined to 2D observation inputs and lack precise perception and comprehension of the 3D physical world — where humans instinctively construct rich, structured mental representations of space, effortlessly aligning objects within a canonical, intuitive, and even personally tailored workspace for manipulation [20, 40, 52, 63, 67]. Therefore, an essential question for the field now is how to effectively equip the VLA models with a profound spatial understanding of the 3D physical world?
@@ II Related Work
3D Foundation Models for Robotics. Some researches [73, 10, 19, 24, 53, 25, 69] have focused on extending the generalist ability of LLMs and VLMs from language-vision towards the 3D world. 3D-LLM [24] integrates a 3D feature extractor with 2D VLMs backbone and train 3D-LLMs on a wide variety of tasks, including dense captioning, 3D question answering, task decomposition, 3D grounding, 3D-assisted dialog, navigation, and so on. LLaVA-3D [10] extends the 2D LLaVA’s capabilities with the proposed 3D patches to bridge 2D features within a 3D space for 3D spatial understanding. Similarly, LEO [25] trains an embodied multi-modal generalist agent that can take egocentric 2D images, 3D point clouds, and texts as task input and handle comprehensive tasks within the 3D environment. Moreover, 3D-VLA [69] builds a generative world model on top of 3D-based LLM to perform 3D reasoning and localization, multimodal goal generation, and embodied action planning. LEO and 3D-VLA are closely related to our work, but their attention is on 3D world understanding and prediction, ignoring the 3D spatial characteristics in the robot action space.
@@ III Methodology
In this section, we describe SpatialVLA model and its training framework in detail. Our model with the proposed Ego3D position encoding and adaptive action grids to capture and learn 3D spatial knowledge for generalizable robot control, which we describe in Sec. III-A. Next, we detail the training procedure of SpatialVLA that consists of a pre-training stage and a post-training stage in Sec. III-B. The pre-training aims to learn generalizable knowledge with large-scale cross-robot data and the goal of post-training is to adapt pre-trained model to specific downstream robot embodiments and tasks.
@@ III-A The SpatialVLA Model Architecture
As illustrated in Fig. 1, SpatialVLA is developed based on a vision-language model to inherit the general world knowledge. Formally, SpatialVLA takes image observations $\mathbf{o}_{t}=\{\mathbf{I}_{t}^{1},..,\mathbf{I}_{t}^{n}\}$ and a natural language task instruction $\mathbf{L}$ as inputs, and then learns a mapping function $\tau(\cdot)$ to generate a sequence of robot actions $\mathbf{A}_{t}=[\mathbf{a}_{t},\mathbf{a}_{t+1},...,\mathbf{a}_{t+H-1}]$, i.e., $\mathbf{A}_{t}=\mathcal{F}(\mathbf{o}_{t},L)$. To empower SpatialVLA with 3D spatial intelligence, we augment the VLM backbone with robotics-specific 3D-aware inputs and outputs, namely, Ego3D Position Encoding and Adaptive Action Grids. The ego3D position encoding representation $\mathbf{O_{3d}}$ aims to capture 3D scene structure via integrating 3D spatial information with 2D semantic features. The adaptive action grids are designed to represent the continuous distribution of robot actions $\mathbf{a}$ with a set of discrete spatial action tokens $\mathfrak{a}=\{\mathfrak{a}^{1},...,\mathfrak{a}^{V}\}$. During training, SpatialVLA model is trained to take the ego3D position encoding representation $\mathbf{O_{3d}}$ and natural language task instruction $\mathbf{L}$ as inputs, and autoregressively generate spatial action tokens $\tilde{\mathfrak{a}}_{t}$ using the cross-entropy objective $\mathcal{L}$,
@@
where the predicted action tokens $\tilde{\mathfrak{a}}_{t}=\tau(\mathbf{O_{3d}},\mathbf{L})$ is the de-tokenized into continuous action signals $\mathbf{a}_{t}$ for robot control. More details of the model architecture and action encoding can be found in Sec. -B.
Ego3D Position Encoding. The proposed Ego3D position encoding integrates depth information from the camera frame and image pixels to construct an egocentric 3D coordinate system, which eliminates the need for robot-camera extrinsic calibration and is agnostic to specific robot setups. Specifically, we use ZoeDepth [4] to estimate depth map $D$ and obtain pixel’s 3D position $\mathbf{p}=\{x,y,z\}$ in the egocentric 3D coordinate system via back-projection $\pi^{-1}$ with camera intrinsic parameters. Then, as illustrated in Fig. 1, we first employ SigLIP [68] visual encoder to extract 2D semantic visual features $\mathbf{X}\in\mathbb{R}^{d\times h\times w}$ to inherit the alignment between vision and language, and calculate their corresponding 3D positions $\mathbf{P}\in\mathbb{R}^{3\times h\times w}$ in the egocentric 3D coordinate system. The egocentric 3D positions $\mathbf{P}$ are then encoded into 3D position embeddings $\mathbf{P}^{{}^{\prime}}\in\mathbb{R}^{d\times h\times w}$ through a sinusoidal function $\boldsymbol{\gamma}(\cdot)$ following by a learnable MLP. The egocentric 3D spatial representations $\mathbf{O_{3d}}\in\mathbb{R}^{d\times h\times w}$ are obtained by adding 3D position embedding $\mathbf{P}^{{}^{\prime}}$ and 2D path visual tokens $\mathbf{X}$, depicted as follows,
@@
As illustrated in Fig. 2, for tokenizing continuous translation and rotation movements, we first normalize each action variable into $[-1,1]$ for each robot environment and statistic the translation and rotation movements $\Delta\mathbf{R}=\{\text{roll, pitch, yaw}\},\;\Delta\mathbf{T}=\{\text{$\phi,\theta,r$}\}$ on the whole dataset mixture (see Sec. -G), following with a parameterized Gaussian distribution fitting $\mathcal{N}(\mu^{a},\Sigma^{a})$. Then, the continuous actions are split into M intervals $\mathbf{G}_{i=1,..,\text{M}}=\{[a_{1}=-1,a_{2}),...,[a_{\text{M-1}},a_{\text{M}}=1]\}$ with equal probability $1/\text{M}$ on each normalized action variable, i.e.,
@@
where $\mathbf{E}_{\text{trans}}\in\mathbb{R}^{d\times\text{M}_{\text{trans}}}$,   $\mathbf{E}_{\text{rot}}\in\mathbb{R}^{d\times\text{M}_{\text{rot}}}$,   $\mathbf{E}_{\text{grip}}\in\mathbb{R}^{d\times 2}$ denote the translation, rotation, and gripper actions, and the total number of action tokens is $V=\text{M}_{\text{trans}}+\text{M}_{\text{rot}}+2$. After training, these learned spatial action tokens capture general robot action knowledge and show a surprising ability in new robot embodiment adaption, as discussed in Sec. III-B. Moreover, it is worth noting that the model only needs to generate 3 tokens for one-step robot actions rather than 7 tokens as in RT-1 [6], RT-2 [7] and OpenVLA [30], achieving in fast model inference speed.
@@ III-B The Pre-training and Post-training Scheme
Pre-training Procedure. We train SpatialVLA from Paligemma2 backbone [62] on a cross-robot dataset mixture with 1.1 Million real robot demonstrations $\{\zeta_{1},...,\zeta_{n}\}$, covering a diverse range of robot embodiments, scenes, and tasks. This pre-training dataset mixture consists of a subset of OXE [13] and the RH20T dataset [18] and we modify the mixture weights from OpenVLA [30] according to the real-word testing performance of individual dataset, which are exhibited in Sec. -A. At the beginning of pre-training, the embeddings $\mathbf{E}_{\mathfrak{a}}$ of spatial action tokens and parameters of MLP in egocentric 3D spatial representation are randomly initialized, and then they are optimized during training, as well as the parameters of vision encoder and LLM backbone. At each training step, a batch of data pairs is extracted at random timesteps $t_{1},...,t_{B}$ from shuffled demonstrations $\{\zeta_{i},...,\zeta_{j}\}$, i.e., a batch of tuple $[(\mathbf{o}_{t_{1}},\mathbf{A}_{t_{1}},\mathbf{L}_{i}),...,(\mathbf{o}_{t_{B}},\mathbf{A}_{t_{B}},\mathbf{L}_{j})]$, and then SpatialVLA is trained with a standard auto-regressive next-token prediction objective in eq. 1. Importantly, the embeddings of text tokens $\mathbf{E}_{\text{text}}$ are frozen to maintain the general world knowledge in pre-trained VLM, and the experimental results show this frozen operation is beneficial for the instruction following ability. Moreover, as discussed in OpenVLA [30], DROID dataset [29] are removed from the data mixture for the final third of pre-training to improve the quality of the pre-trained SpatialVLA model.
@@ IV Experiment
To answer these questions, as shown in Fig. 3, we evaluate SpatialVLA’s capabilities across a representative spectrum of 7 different robot learning scenarios with 24 real-robot tasks and 3 simulation environments. Firstly, we evaluate SpatialVLA in both SimplerEnv [35] simulation and the real-world WidowX robot platform (BridgeV2 [64] [64] setups), testing its out-of-the-box control capabilities on different robots with setups matching the pre-training dataset. Second, we assess the fine-tuning efficacy of our method in both simulation and real-world settings, including LIBERO [36] and new Franka robot setups, to adapt to new robot environments and tasks. Then, we design 4 special tasks that require precise spatial understanding in 2 different real-world robot environments to test the effectiveness of spatial representations of SpatialVLA. Finally, we conduct comprehensive ablation studies on a mixture of Fractal [6] and BridgeV2 [64] datasets to verify our design decisions in SpatialVLA. For more details on evaluation setups, see Sec. -E.
Implementation Details. The SpatialVLA model is pre-trained with 1.1 Million real-robot demonstrations from the OXE  [13] and RH20T dataset [18] on a cluster of 64 A100 GPUs for 10 days, using a batch size of 2048. For input robot observation, the SpatialVLA policy is only conditioned on one third-person camera and takes one image for constructing egocentric 3D spatial representations. For output robot actions, the SpatialVLA policy predicts a chunk of $T=4$ future actions (12 spatial action tokens from total $V=8194$ tokens) and executes the ensemble actions before predicting the next chunk. During inference, SpatialVLA requires 8.5GB of GPU memory and runs at approximately 20Hz on one NVIDIA RTX 4090 GPU to run evaluations in both simulation and real-world. For more details about model training and deployment, please refer to the Sec. -D.
@@ IV-A Performing Zero-shot Robot Control
For a more comprehensive evaluation, we conduct experiments on a real-world WidowX robot platform from the BridgeData V2 evaluation [64]. As shown in Fig. 4, we design seven task suites for the WidowX robot, encompassing language grounding, semantic understanding (unseen background and poses), and motion distractors (manually move the object). All generalist manipulation policies, including Octo, RT-1-X, OpenVLA, and RoboVLM, are evaluated across 7 task suites with 11 trials each, resulting in a total of 77 rollouts. A more detailed breakdown of all tasks and policy settings can be found in the Sec. -E.
SimplerEnv Evaluation of Google Robot and WidowX. Tab. I summarizes the zero-shot and fine-tuning results across different manipulation policies on the Google Robot setup. On average, SpatialVLA achieves the highest overall visual matching and variant aggregation performance with a significant margin. Our SpatialVLA model yields 71.9% and 75.1% Visual Matching scores in zero-shot and fine-tuning settings, surpassing the second-best policy, RoboVLM, by +15.6% and +11.7% margins. Notably, our model trained from scratch on OXE mixture with RH20T surpasses the state-of-the-art closed-source model RT-2-X [13], achieving superior performance in Visual Matching (71.9% vs 60.7%) and Variant Aggregation (68.8% vs 64.3%), while using significantly fewer model parameters (3.5B vs 55B). Qualitatively, we find that SpatialVLA exhibits greater generalizability and robustness across diverse robotic manipulation tasks and environmental conditions, characterized by varying visual appearances, which is further supported by its superior performance in variant aggregation. In particular, SpatialVLA also matches or outperforms the latest SOTA model $\pi_{0}$. Tab. II summarizes the results across different manipulation policies on the WidowX setup. Our model surpasses the state-of-the-art RoboVLM policy, achieving overall success rates of 34.4% and 42.7%. Fine-tuning on the BridgeV2 yields a remarkable 100% success rate in the “Put Eggplant in Yellow Basket” task, demonstrating the model’s exceptional zero-shot manipulation capability.
@@
Real-world WidowX Evaluation. Fig. 4 presents the results of the real-world “out-of-the-box” evaluation in WidowX robot platform. We observe that, in simple single-task scenarios (#1 close microwave), all the policies exhibit some generalizability, successfully completing tasks in unseen environments. However, in moderately complex tasks (#3-7), most policies, such as RT-1-X, Octo, and RoboVLM struggle with manipulation, frequently encountering issues like object misidentification and grasp failures. Compared to OpenVLA, our method demonstrates superior robustness in handling motion disturbances (human-induced dynamic object movement in tasks #3 and #4), successfully tracking and grasping carrot and eggplant. Furthermore, in the instruction-following tasks (#5-7), our method demonstrates strong instruction-following ability, accurately executing tasks like picking up a green cup and placing it on a white plate, not a pink one, based on color descriptions in the prompts, outperforming OpenVLA and other generalist policies. Overall, SpatialVLA achieves a higher average success rate, showcasing robust and generalizable operation capabilities in unseen scenarios, objects, language grounding, and dynamic motions.